OpenStudy (anonymous):

Hey I need some help with functions, I have it most of the way I just wanted to know what I have to do after this.

OpenStudy (anonymous):

Let f(x) = x^2 + x + 2 and g(x) = 2x^2 + 5. Find f(g(x)). Show each step of your work. @amistre64

OpenStudy (anonymous):

all I do is substitute g(x) everywhere there is an x in the equation right? f(2x^2+5)=(2x^2+5)^2+2x^2+5)+2 then simplify to f(sx^2+5)=4x^4+25)+(2x^2+5)+2

OpenStudy (anonymous):

no I messed up on the beginning; then simplify to f(2x^2+5)=(4x^4+25)+2x^2+5)+2

OpenStudy (anonymous):

@Directrix would you know how to do this?

Directrix (directrix):

I'll have to begin my work from the top. Hold on.

OpenStudy (anonymous):

ok thank you, ill wait patiently :)

Directrix (directrix):

Is this how you set up your solution (so far)?